A piece of land on the Griswold Scout Reservation has been clear-cut, and the property owner is deciding whether to replant. It costs $20,000 to do the initial tree replanting and an additional maintenance cost of $5,000 both one year later and two years later. Alternatively, GSR can earn 3% on its investments. The optimal harvest time for the planted trees is known to be 40 years of growth. At that time, GSR will receive $100,000.

a) Calculate the profitability of replanting.

b) GSR’s deed requires this piece of land to have a conservation easement within 50 years. If they replant, they will still receive harvest revenue as in part a), and now will also incur $5,000 of costs for site preparation 5 years after harvesting and then 5 years after that receive $10,000 payment from the Lakes Region Conservation Trust for granting the easement. Alternatively, GSR can choose not to replant now, incur the site prep costs 5 years from today and receive the easement revenue from LRCT 5 years after that. Show the profitability of each alternative.
Answer #1

Assuming profitability is return on investment in current terms.

The cashflows and NPV, based on a 3% rate, for the cashflows is shown below.

Year Cashflow
0 -20000
1 -5000
2 -5000
3 0
4 0
5 0
6 0
7 0
8 0
9 0
10 0
11 0
12 0
13 0
14 0
15 0
16 0
17 0
18 0
19 0
20 0
21 0
22 0
23 0
24 0
25 0
26 0
27 0
28 0
29 0
30 0
31 0
32 0
33 0
34 0
35 0
36 0
37 0
38 0
39 0
40 100000
NPV $1,088

The NPV has been calculated by the formula =NPV(rate,cashflow). Alternatively, it can be manually calculated by

NPV=-20000-5000/1.03-5000/1.032+100000/1.0340

Both will give the same answer.

Profitability=return/investment=1088/20000=5.44%

B. If they replant, the cashflow would be

Year Cashflow
0 -20000
1 -5000
2 -5000
3 0
4 0
5 0
6 0
7 0
8 0
9 0
10 0
11 0
12 0
13 0
14 0
15 0
16 0
17 0
18 0
19 0
20 0
21 0
22 0
23 0
24 0
25 0
26 0
27 0
28 0
29 0
30 0
31 0
32 0
33 0
34 0
35 0
36 0
37 0
38 0
39 0
40 100000
41 0
42 0
43 0
44 0
45 -5000
46 0
47 0
48 0
49 0
50 10000
NPV $2,047

Since NPV is 2047, profitability=2047/20000=10.235%

If they do not replant, the cashflow would be

Year Cashflow
0 0
1 0
2 0
3 0
4 0
5 -5000
6 0
7 0
8 0
9 0
10 10000
NPV 3127.90

But since the investment is after 5 years, to calculate profitability we also need its NPV today. NPV of 5000 after 5 years is

NPV=-5000/1.035

=4313

Hence, profitability=3127.90/4313=72.52%